# ModeTv2: GPU-accelerated Motion Decomposition Transformer for Pairwise Optimization in Medical Image Registration

### v1

@@ Methods
HD95 (mm)
ASSD (mm)
%$\%|J_{\phi}|\leq 0$
Initial
34.0$\pm$1.9
6.62$\pm$0.44
2.04$\pm$0.13
SyN (Avants et al., 2008)
53.0$\pm$2.0
5.65$\pm$0.52
1.44$\pm$0.13
$<$0.00006%
VM (Balakrishnan et al., 2019)
54.7$\pm$1.7
6.07$\pm$0.47
1.50$\pm$0.12
$<$0.2%
VM-diff (Dalca et al., 2019)
51.2$\pm$2.5
6.02$\pm$0.46
1.54$\pm$0.12
$<$0.004%
TM (Chen et al., 2022a)
58.5$\pm$1.7
5.83$\pm$0.50
1.38$\pm$0.13
$<$0.2%
TransMatch (Chen et al., 2024)
56.0$\pm$1.7
5.84$\pm$0.50
1.42$\pm$0.13
$<$0.09%
PR++ (Kang et al., 2022)
58.4$\pm$1.6
5.85$\pm$0.48
1.39$\pm$0.12
$<$0.08%
DMR (Chen et al., 2022b)
58.1$\pm$1.7
5.76$\pm$0.51
1.38$\pm$0.13
$<$0.09%
PCnet (Lv et al., 2022)
61.0$\pm$1.7
5.59$\pm$0.52
1.30$\pm$0.13
$<$0.003%
PIViT (Ma et al., 2023):
56.4$\pm$1.6
5.48$\pm$0.50
1.35$\pm$0.13
$<$0.02%
I2G (Liu et al., 2022)
57.4$\pm$1.5
5.55$\pm$0.44
1.34$\pm$0.11
$<$0.005%
ModeT (Wang et al., 2023)
60.2$\pm$1.6
5.38$\pm$0.46
1.27$\pm$0.12
$<$0.007%
ModeTv2-s
61.5$\pm$1.6
5.29$\pm$0.49
1.23$\pm$0.12
$<$0.005%
ModeTv2-diff-s
61.1$\pm$1.6
5.30$\pm$0.48
1.24$\pm$0.12
$<$0.0008%
ModeTv2-l
62.3$\pm$1.6
5.21$\pm$0.49
1.21$\pm$0.12
$<$0.02%
ModeTv2-diff-l 62.8$\pm$1.7 5.11$\pm$0.52 1.17$\pm$0.13 $<$0.002%
@@
Figure 7: The visualization of the registration results and the corresponding deformation fields from different methods on three datasets.
@@
Figure 9: Illustration of the DSC results of different DL methods on three datasets during the first ten epochs of the training phase.
The numerical results of different methods on datasets ABCT, LPBA and Mindboggle are listed in Tables 12 and 3, respectively. It can be observed that our ModeTv2, especially the large model, consistently attained satisfactory registration accuracy with respect to the metrics of DSC, HD95 and ASSD. These Tables also report the percentage of voxels with non-positive Jacobian determinant (%$\%|J_{\phi}|\leq 0$), which demonstrate our ModeTv2 predicted high quality non-rigid deformation field. We now analyze the comparison results on each dataset. ABCT. It can be seen from Table 1 that even with affine pre-alignment, the ABCT images still had large deformation (initial DSC value of 45.7%). In such case, traditional SyN and single-stage networks did not perform well. In contrast, our ModeTv2 achieved overall the best results of DSC, HD95 and ASSD. In particular, the large version of ModeTv2 outperformed all comparison methods (even including its small counterpart) obviously. Moreover, the diffeomorphism version of ModeTv2-s generated the lowest negative Jacobian ratio, while maintaining satisfactory registration accuracy. LPBA. As shown in Table 2, the small version of ModeTv2 already generated the best registration accuracy among all other comparison methods, and the large version further improved the performance. Compared with the normal version, the diffeomorphism version of ModeTv2-l achieved almost the same registration accuracy and maintained extremely low negative Jacobian ratio. Mindboggle. The images from Mindboggle possess complex deformation and therefore the initial affine pre-alignment attained only an average DSC of 34.0%, as listed in Table 3. Among all methods, our ModeTv2 again achieved favorable registration accuracy and generated high quality deformation field. It is worth noting that for all the experimental datasets, our method in this study consistently outperformed the conference version ModeT (Wang et al., 2023), which demonstrates the efficacy of our simple RegHead in fusion of multiple deformation subfields. Moreover, due to the CUDA implementation, the ModeTv2 is much more time-/memory-efficient compared with ModeT, which will be discussed later.
The distributions of DSC results of specific regions on datasets ABCT, LPBA and Mindboggle are shown in Figs. 45 and 6, respectively. In particular, Fig. 4 illustrates the DSC distributions of four organs including liver, spleen, left kidney and right kidney; Fig. 5 illustrates the DSC distributions of seven anatomical regions including frontal, parietal, occipital, temporal, cingulate, hippocampus and putamen; Fig. 6 illustrates the DSC distributions of five anatomical regions including frontal, parietal, occipital, temporal and cingulate. It can be observed that for most anatomical regions, the DSC distributions of our large ModeTv2 were positioned higher to some perceptible extent.
Figs. 7 visualizes the registered images and the corresponding deformation fields from different methods on three datasets (we encourage you to zoom in for better visualization). Our ModeTv2 generated more accurate registered images, and the internal structures can be consistently preserved by using our method. Moreover, our estimated deformation fields were relatively smooth, even with large displacements, the local deformations still seem reasonable. Fig. 8 takes the registration of one image pair as an example to show the multi-level deformation fields generated by our method using ModeTv2-diff-s. Our ModeTv2 effectively modeled multiple motion modes and our RegHead fused them together. The final deformation field $\phi$ accurately warped the moving image to registered with the fixed image.
@@ 4.5.4 Registration Efficiency
Table 4: The time and memory usage of different methods during trainning and inference .
Methods
Training Time ($ms$)
Training GPU Mem($MB$)
Inference Time ($ms$)
Inference GPU Mem($MB$)
SyN (Avants et al., 2008)
[…]
VM (Balakrishnan et al., 2019)
1078.94
9174
102.5
3677
VM-diff (Dalca et al., 2019)
283.73
4120
27.43
1974
TM (Chen et al., 2022a)
1243.14
17906
201.7
5263
TransMatch (Chen et al., 2024)
1225.95
22922
240.85
4392
PR++ (Kang et al., 2022)
3111.99
17584
522.7
9435
DMR (Chen et al., 2022b)
3112.99
22979
980.4
8383
PCnet (Lv et al., 2022)
3436.53
23584
753.87
6324
PIViT (Ma et al., 2023):
654.05
4506
28.24
2085
I2G (Liu et al., 2022)
1195.74
19558
290.2
9805
ModeT (Wang et al., 2023)
7595.52
18710
558.3
9809
ModeTv2-s
1069.74
8960
235.07
3530
ModeTv2-diff-s
1166.65
9520
258.83
4098
ModeTv2-l
1582.76
19808
277.31
6328
ModeTv2-diff-l
1696.38
20450
301.77
6408
@@ Methods
By observing Figure 11, it is evident that VM and VM-diff had the smallest parameter counts, while the Transformer-based TM and TransMatch largely increased the number of parameters. Our proposed small ModeTv2 had a relatively smaller number of parameters, approaching the magnitude of VM, whereas the large ModeTv2’s parameter count was still smaller than TM and TransMatch.
@@ 5.1 Discussion on Registration Accuracy
According to Tables 1, 2, and 3, our ModeTv2 demonstrates a high level of accuracy in registration. The results across all three datasets surpass those of our conference version of ModeT, by +1.1%, +0.9%, +1.3% on ABCT, LPBA, and Mindboggle datasets, respectively. This improvement is mainly attributed to the design of RegHead. RegHead serves two purposes. One is to integrate multiple motion patterns and reevaluate unreasonable neighborhood relationships. Additionally, with only one convolutional layer, RegHead minimally transforms semantic information for multiple motion subfields, maintaining the explicit semantic generation of the ModeTv2 operator. This facilitates transferability, allowing for fine-tuning of the encoder during PO without relearning the deformation field estimation.
@@ 6 Conclusion and Future Work
We present a GPU-accelerated motion decomposition Transformer, namely ModeTv2, to naturally model the correspondence between images and convert it into deformation fields, which improves the interpretability and efficiency of the deep-learning-based registration networks. The proposed ModeTv2 is built upon our MICCAI work ModeT. It modifies the fusion method of multiple subfields from the original competitive weighting to the current RegHead. This solves the problem that subfields in ModeT take the neighborhood relationship into account. In order to provide efficient computing, which is an important issue for the volumetric registration, we design CUDA operators in ModeTv2. Importantly, for different scenarios, we provide four versions of ModeTv2, i.e., the small model, large model, and their corresponding diffeomorphism models. Experimental results demonstrate that the proposed method achieves superior performance over existing techniques in terms of both registration accuracy and efficiency. We also conduct pairwise optimization experiments across three datasets. By adopting the PO, our ModeTv2 balances accuracy, efficiency, and generalizability.
@@ Acknowledgments
This work was supported in part by the National Natural Science Foundation of China under Grants 62071305, 62171290 and 12326619, in part by the Guangdong-Hong Kong Joint Funding for Technology and Innovation under Grant 2023A0505010021, in part by the Guangdong Basic and Applied Basic Research Foundation under Grant 2022A1515011241, and in part by the Shenzhen Science and Technology Program under Grant SGDX20201103095613036.

### v2

@@ 4.5.1 Registration Accuracy
Initial 34.0$\pm$1.9 6.62$\pm$0.44 2.04$\pm$0.13 -
SyN [2] 53.0$\pm$2.0 5.65$\pm$0.52 1.44$\pm$0.13 $<$0.000006%
VM [6] 54.7$\pm$1.7 6.07$\pm$0.47 1.50$\pm$0.12 $<$1%
VM-diff [16] 51.2$\pm$2.5 6.02$\pm$0.46 1.54$\pm$0.12 $<$0.009%
TM [9] 58.5$\pm$1.7 5.83$\pm$0.50 1.38$\pm$0.13 $<$0.9%
TransMatch [14] 56.0$\pm$1.7 5.84$\pm$0.50 1.42$\pm$0.13 $<$1%
PR++ [35] 58.4$\pm$1.6 5.85$\pm$0.48 1.39$\pm$0.12 $<$0.5%
DeFormer [12] 58.1$\pm$1.7 5.76$\pm$0.51 1.38$\pm$0.13 $<$0.7%
PCnet [44] 61.0$\pm$1.7 5.59$\pm$0.52 1.30$\pm$0.13 $<$0.006%
PIViT [45]: 56.4$\pm$1.6 5.48$\pm$0.50 1.35$\pm$0.13 $<$0.05%
I2G [42] 57.4$\pm$1.5 5.55$\pm$0.44 1.34$\pm$0.11 $<$0.02%
ModeT [69] 60.2$\pm$1.6 5.38$\pm$0.46 1.27$\pm$0.12 $<$0.03%
ModeTv2-s 61.5$\pm$1.6 5.29$\pm$0.49 1.23$\pm$0.12 $<$0.02%
ModeTv2-diff-s 61.1$\pm$1.6 5.30$\pm$0.48 1.24$\pm$0.12 $<$0.0002%
ModeTv2-l 62.3$\pm$1.6 5.21$\pm$0.49 1.21$\pm$0.12 $<$0.08%
ModeTv2-diff-l 62.8$\pm$1.7 5.11$\pm$0.52 1.17$\pm$0.13 $<$0.007%
@@
Table 4: The numerical results of different registration methods on the IXI (30 ROIs) dataset.
[…]
Initial 40.6$\pm$3.5 6.24$\pm$0.63 2.72$\pm$0.29 -
SyN [2] 63.6$\pm$3.9 5.55$\pm$0.88 1.74$\pm$0.28 $<$0.0002%
VM [6] 72.4$\pm$2.5 2.90$\pm$0.36 1.08$\pm$0.12 $<$0.05%
VM-diff [16] 57.5$\pm$3.6 5.11$\pm$0.85 1.85$\pm$0.25 $<$0.003%
TM [9] 73.9$\pm$2.7 2.82$\pm$0.37 1.03$\pm$0.13 $<$0.2%
TransMatch [14] 73.6$\pm$2.6 2.89$\pm$0.39 1.05$\pm$0.13 $<$0.2%
PR++ [35] 74.4$\pm$2.4 2.78$\pm$0.37 1.01$\pm$0.12 $<$0.04%
DeFormer [12] 71.5$\pm$3.2 3.08$\pm$0.44 1.14$\pm$0.15 $<$0.05%
PCnet [44] 74.9$\pm$2.5 2.77$\pm$0.39 0.99$\pm$0.12 $<$0.00006%
PIViT [45] 75.2$\pm$2.1 2.68$\pm$0.37 0.97$\pm$0.10 $<$0.03%
I2G [42] 74.2$\pm$2.4 2.80$\pm$0.37 1.01$\pm$0.12 $<$0.0003%
ModeT [69] 75.0$\pm$2.2 2.74$\pm$0.37 0.98$\pm$0.11 $<$0.0002%
ModeTv2-s 75.8$\pm$2.2 2.68$\pm$0.37 0.96$\pm$0.11 $<$0.0004%
ModeTv2-diff-s 75.7$\pm$2.2 2.67$\pm$0.37 0.96$\pm$0.11 0
ModeTv2-l 75.9$\pm$2.2 2.64$\pm$0.37 0.95$\pm$0.11 $<$0.02%
ModeTv2-diff-l 75.7$\pm$2.3 2.69$\pm$0.39 0.96$\pm$0.11 0
Figure 7: Box plots showing the distributions of DSC scores of seven regions on the IXI dataset produced by different registration methods.
The numerical results of different methods on datasets ABCT, LPBA, Mindboggle and IXI are listed in Tables 123 and 4, respectively. It can be observed that our ModeTv2, especially the large model, consistently attained satisfactory registration accuracy with respect to the metrics of DSC, HD95 and ASSD. These Tables also report the percentage of voxels with non-positive Jacobian determinant (%$\%|J_{\phi}|\leq 0$), which demonstrate our ModeTv2 predicted high quality non-rigid deformation field. We now analyze the comparison results on each dataset. ABCT. It can be seen from Table 1 that even with affine pre-alignment, the ABCT images still had large deformation (initial DSC value of 45.7%). In such case, traditional SyN and single-stage networks did not perform well. In contrast, our ModeTv2 achieved overall the best results of DSC, HD95 and ASSD. In particular, the large version of ModeTv2 outperformed all comparison methods (even including its small counterpart) obviously. Moreover, the diffeomorphism version of ModeTv2-s generated the lowest negative Jacobian ratio, while maintaining satisfactory registration accuracy. LPBA. As shown in Table 2, the small version of ModeTv2 already generated the best registration accuracy among all other comparison methods, and the large version further improved the performance. Compared with the normal version, the diffeomorphism version of ModeTv2-l achieved almost the same registration accuracy and maintained extremely low negative Jacobian ratio. Mindboggle. The images from Mindboggle possess complex deformation and therefore the initial affine pre-alignment attained only an average DSC of 34.0%, as listed in Table 3. Among all methods, our ModeTv2 again achieved favorable registration accuracy and generated high quality deformation field. IXI. The IXI dataset differs from LPBA and Mindboggle as its labeled ROIs are not predominantly concentrated in the lobar regions but are instead located near the brainstem. Consequently, this dataset evaluates overall alignment performance rather than intricate deformation, leading to higher DSC scores for most methods. As illustrated in Tables 3 and 4, PIViT, a lightweight pyramid network leveraging the Swin Transformer structure, achieved a DSC of 75.2% on IXI, despite its moderate performance on Mindboggle. In contrast, our model achieved the highest DSC score, with the -diff models further ensuring fold-free deformations. It is worth noting that for all the experimental datasets, our method in this study consistently outperformed the conference version ModeT [69], which demonstrates the efficacy of our simple RegHead in fusion of multiple deformation subfields. Moreover, due to the CUDA implementation, the ModeTv2 is much more time-/memory-efficient compared with ModeT, which will be discussed later.
The distributions of DSC results of specific regions on datasets ABCT, LPBA, Mindboggle and IXI are shown in Figs. 456 and 7, respectively. In particular, Fig. 4 illustrates the DSC distributions of four organs including liver, spleen, left kidney and right kidney; Fig. 5 illustrates the DSC distributions of seven anatomical regions including frontal, parietal, occipital, temporal, cingulate, hippocampus and putamen; Fig. 6 illustrates the DSC distributions of five anatomical regions including frontal, parietal, occipital, temporal and cingulate; Fig. 7 presents the DSC distributions across seven anatomical regions: white matter, cortex, ventricle, caudate, putamen, cerebrospinal fluid (CSF), and hippocampus. It can be observed that for most anatomical regions, the DSC distributions of our large ModeTv2 were positioned higher to some perceptible extent.
Figure 8: The visualization of the registration results and the corresponding deformation fields from different methods on four datasets.
@@
Fig. 8 visualizes the registered images and the corresponding deformation fields from different methods on four datasets (we encourage you to zoom in for better visualization). Our ModeTv2 generated more accurate registered images, and the internal structures can be consistently preserved by using our method. Moreover, our estimated deformation fields were relatively smooth, even with large displacements, the local deformations still seem reasonable. Fig. 9 takes the registration of one image pair as an example to show the multi-level deformation fields generated by our method using ModeTv2-diff-s. Our ModeTv2 effectively modeled multiple motion modes and our RegHead fused them together. The final deformation field $\phi$ accurately warped the moving image to registered with the fixed image.
Figure 10: Illustration of the DSC results of different DL methods on four datasets during the first ten epochs of the training phase.
@@ 4.5.4 Registration Efficiency
Table 5: The time and memory usage of different methods during training and inference.
Methods Training Time ($ms$) Training GPU Mem($MB$) Inference Time ($ms$) Inference GPU Mem($MB$)
SyN [2] - - $>$105 -
VM [6] 1078.94 9174 102.5 3677
VM-diff [16] 283.73 4120 27.43 1974
TM [9] 1243.14 17906 201.7 5263
TransMatch [14] 1225.95 22922 240.85 4392
PR++ [35] 3111.99 17584 522.7 9435
DeFormer [12] 3112.99 22979 980.4 8383
PCnet [44] 3436.53 23584 753.87 6324
PIViT [45]: 654.05 4506 28.24 2085
I2G [42] 1195.74 19558 290.2 9805
ModeT [69] 7595.52 18710 558.3 9809
ModeTv2-s 1069.74 8960 235.07 3530
ModeTv2-diff-s 1166.65 9520 258.83 4098
ModeTv2-l 1582.76 19808 277.31 6328
ModeTv2-diff-l 1696.38 20450 301.77 6408
@@
By observing Fig. 12, it is evident that VM and VM-diff had the smallest parameter counts, while the Transformer-based TM and TransMatch largely increased the number of parameters. Our proposed small ModeTv2 had a relatively smaller number of parameters, approaching the magnitude of VM, whereas the large ModeTv2’s parameter count was still smaller than TM and TransMatch.
[…]
@@ 5.1 Discussion on Registration Accuracy
According to Tables 123 and 4, our ModeTv2 demonstrates a high level of accuracy in registration. The results across all four datasets surpass those of our conference version of ModeT, by +1.1%, +0.9%, +1.3%, +0.8% on ABCT, LPBA, Mindboggle, and IXI datasets, respectively. This improvement is mainly attributed to the design of RegHead. RegHead serves two purposes. One is to integrate multiple motion patterns and reevaluate unreasonable neighborhood relationships. Additionally, with only one convolutional layer, RegHead minimally transforms semantic information for multiple motion subfields, maintaining the explicit semantic generation of the ModeTv2 operator. This facilitates transferability, allowing for fine-tuning of the encoder during PO without relearning the deformation field estimation.
@@ 5.3 Discussion on PO Ability
5.4 Discussion on Interpretability
Currently, deep learning methods for image registration lack sufficient interpretability. In the case of same-domain tasks, enhancing the interpretability of DL registration models can improve the plausibility of the resulting deformation fields and accelerate model convergence. Firstly, by enhancing interpretability, we enable the model to compute deformation fields in a pre-defined, reasonable manner. This contrasts with other DL-based registration methods that only impose constraints on the output, thus enhancing the plausibility of the final deformation fields. Secondly, increasing interpretability effectively strengthens the model’s inductive bias for registration, which, in turn, accelerates convergence.
As we mentioned in the Introduction Section, we argue that adopting network designs closely aligned with the computational aspects of the registration task can alleviate the learning burden on the network and enhance convergence efficiency. Some methods use cross-attention mechanisms to capture spatial relationships between moving and fixed image features, aiming to enhance the model’s inductive bias for registration. However, most current methods utilize this mechanism merely to enhance feature representation, with few considering its direct usage for solving the deformation field. In contrast, our motion decomposition transformer simultaneously considers the decomposition of motion patterns in features and explicitly calculates the deformation field from the correlations of features, showcasing interpretability.
Regarding the plausibility of the deformation field, we observe from our accuracy experiments that the ModeT series achieve high accuracy while maintaining a low folding rate. Even without using a diffeomorphic approach, the deformation fields generated by ModeTv2 are already quite plausible, and the deformation fields produced by ModeTv2-diff exhibit the lowest folding rate among DL methods. Regarding convergence, our experimental results show that the ModeTv2 approaches convergence after just the first iteration. This is due to our pre-defined approach for computing deformation fields in the decoder part of the network, which means that our model requires minimal learning on how to generate the deformation field. Instead, it focuses more on learning how to extract image features, significantly improving convergence speed.
As for tasks in different domains, our interpretable model demonstrates strong generalization capabilities. During the PO experiments, the network focuses primarily on feature extraction, while the interpretable deformation field computation requires little to no retraining. The PO experiments have demonstrated the strong cross-domain ability of our model.
@@ 6 Conclusion and Future Work
We present a GPU-accelerated motion decomposition Transformer, namely ModeTv2, to naturally model the correspondence between images and convert it into deformation fields, which improves the interpretability and efficiency of the deep-learning-based registration networks. The proposed ModeTv2 is built upon our MICCAI work ModeT. It modifies the fusion method of multiple subfields from the original competitive weighting to the current RegHead. This solves the problem that subfields in ModeT take the neighborhood relationship into account. In order to provide efficient computing, which is an important issue for the volumetric registration, we design CUDA operators in ModeTv2. Importantly, for different scenarios, we provide four versions of ModeTv2, i.e., the small model, large model, and their corresponding diffeomorphism models. Experimental results demonstrate that the proposed method achieves superior performance over existing techniques in terms of both registration accuracy and efficiency. We also conduct pairwise optimization experiments across different datasets. By adopting the PO, our ModeTv2 balances accuracy, efficiency, and generalizability.
@@ Acknowledgments
This work was supported in part by the National Natural Science Foundation of China under Grants 62471306, 62071305, 62171290 and 12326619, in part by the Guangdong-Hong Kong Joint Funding for Technology and Innovation under Grant 2023A0505010021, in part by the Shenzhen Medical Research Fund under Grant D2402010, and in part by the Guangdong Basic and Applied Basic Research Foundation under Grant 2022A1515011241.